DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, July 31, 2003 Thompson 235 2:30 pm (Note: Special Day, Time and Room!)

Integrating factors and non-elliptic Schrodinger equations

Luis Vega (Universidad del Pais Vasco, Spain)

I’ll present a joint work with C. Kenig, G. Ponce and C. Rolvung on the initial value problem of a semilinear Schrodinger equation. The matrix of the second order terms is real, symmetric and just nondegenerate. The key of the proof is the construction of an integrating factor using a non-standard pesudodifferential calculus.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.